# Truth or Twist? Optimal Model Selection for Reliable Label Flipping Evaluation in LLM-based Counterfactuals

### v1

@@ 5 Results
Judge model performance depends on its relationship to the generator. As shown in Table 1 ($\Delta$ is calculated by subtracting the user study results from the results of the judge models $LLM_{\mathcal{J}}$ results), the alignment of $LLM_{\mathcal{J}}$ with human annotators depends heavily on their relationship $\mathcal{R}(\mathcal{G},\mathcal{J})$. To measure the alignment, we (1) count instances where $LLM_{\mathcal{J}}$ most closely or least matched human annotators across three datasets, (2) compute the ratio of most-to-least alignment ($r_{m/l}\uparrow$). $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imwo}$ relationship achieves the highest alignment ($r_{m/l}=2.17$). In contrast, $\mathcal{R}_{dm}$ ($r_{m/l}=0.83$) or $\mathcal{R}_{imw}$ ($r_{m/l}=0.54$) frequently show poor alignment with human annotators potentially due to data contamination Li et al. (2025), as these models are either fine-tuned on the target dataset or share architectural similarities with $LLM_{\mathcal{G}}$. Such overlap could bias $LLM_{\mathcal{J}}$ in its evaluation of label flipping, leading to either overestimating or underestimating the LFR.
@@
Despite $LLM_{\mathcal{J}}$ with $\mathcal{R}_{dm}$ or $\mathcal{R}_{imw}$ relationships achieving the highest downstream task performance across the target datasets (Table 3), their alignment with human annotators in evaluating label flipping remains considerably weak (Table 1). This contrast highlights that strong downstream task performance does not lead to reliable label flipping evaluation, and it underscores the need for careful judge model selection.
@@
LLMs may struggle to capture nuanced changes when determining the label of imperfect counterfactuals, as the context may not have fully shifted to support a different label, resulting in ambiguity (Appendix E). Notably, even the optimal judge model fails to fully capture label flipping, exhibiting an average discrepancy of 22.78% relative to the user study results, which implies that a fully automated CDA pipeline is insufficient and necessitates human oversight.
@@
We notice that when the LFR of $LLM_{\mathcal{J}}$, given its relationship to $LLM_{\mathcal{G}}$, aligns more closely with user study outcomes, the labels identified by $LLM_{\mathcal{J}}$ are associated with improved performance and greater robustness on both unseen and out-of-distribution data Kaushik et al. (2020); Gardner et al. (2020) (Appendix G). This association is particularly evident when these identified labels are treated as the ground-truth labels for counterfactuals, which subsequently serve as data points for CDA. This can be attributed to the fact that noisy and incorrect labels provided by judge models with suboptimal relationships contribute to performance deterioration Zhu et al. (2022); Song et al. (2023).
@@ 6 Conclusion
In our experiments, we exclusively use models from Qwen and Llama families to generate counterfactuals, as from DeepSeek-R1 DeepSeek-AI (2025) distilled Qwen2.5 and Llama3 models are officially provided [^8] https://huggingface.co/collections/deepseek-ai/deepseek-r1-678e1e131c0169c0bc89728d and can be used out-of-the-box. Additional work is required when employing models from a different model family as $LLM_{\mathcal{G}}$, including using DeepSeek-R1 to generate synthetic data and fine-tuning $LLM_{\mathcal{G}}$ to derive $LLM_{\mathcal{J}}$ with distillation relationships.
For label flipping evaluation, we select BERT, RoBERTa, Phi4-14B, and Mistral-Large-Instruct-2411 as representatives of independent3.2) open-source LLMs with different parameter sizes, without comprehensively assessing models from all other widely known model families. In particular, we evaluate only open-source models, rather than closed-source, proprietary models.
@@
Appendix A Datasets
A.1 Label Distribution
@@ Appendix B Models
DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI (2025) 7B https://huggingface.co/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B
@@
Appendix C Downstream Task Performance
@@
Appendix D Human Annotation
@@
Appendix E Challenges in Label Flipping Identification
@@
Appendix F Automatic Evaluation: A Sanity Check

### v2

@@ 5 Results
Figure 2: The average ranking of judge-generator model relationship based on $\Delta$ from Table 4 in Appendix G (lower rankings indicate better alignment). Counterfactual examples are generated by Qwen2.5-{14B,32B} and Llama3-{8B,70B}, evaluated across judge models exhibiting same, distilled, same family, ensemble, and independent w/ and w/o fine-tuning relationships on AG News, SNLI and SST2.
Judge model performance depends on its relationship to the generator.
As shown in Figure 3(a) and Figure 3(b), $LLM_{\mathcal{J}}$ with $\mathcal{R}_{imwo}$ relationship achieves the highest alignment with human annotators ($rank=4.15$, $r_{m/l}=3.5$, $\rho=0.47$). In contrast, $\mathcal{R}_{dm}$ ($rank=5.58$, $r_{m/l}=1$, $\rho=0.38$) or $\mathcal{R}_{imw}$ ($rank=5.86$, $r_{m/l}=0.23$, $\rho=0.29$) demonstrate poor alignment with human judgments. This may attribute to data contamination Li et al. (2025), as these models are either fine-tuned on the target dataset or share architectural similarities with $LLM_{\mathcal{G}}$. Such overlap could bias $LLM_{\mathcal{J}}$ in its evaluation of label flipping, potentially leading to either overestimating or underestimating the LFR. Notably, ensembling results from all judge models does not necessarily lead to better alignment, as it partially relies on results from suboptimal judge models. Additionally, we find the first two measures (average ranking and $r_{m/\ell}$) are moderately correlated, with a Spearman correlation coefficient of 0.67, indicating that $r_{m/\ell}$ also serves as a reliable metrics to capture alignment between model pairs and human judgments.
(a) Average ranking of the relationships $\mathcal{R}_{\mathcal{G},\mathcal{J}}$.
(b) Most-to-least ratio ($r_{m/\ell}$) of the relationships $\mathcal{R}_{\mathcal{G},\mathcal{J}}$.
Figure 3: Average ranking and most-to-least ratio for the relationships $\mathcal{R}_{\mathcal{G},\mathcal{J}}$: same, distilled, same family, ensemble, and independent w/ and w/o fine-tuning.
@@
Despite $LLM_{\mathcal{J}}$ with $\mathcal{R}_{dm}$ or $\mathcal{R}_{imw}$ relationships achieving the highest downstream task performance across the target datasets (Table 2 in Appendix D), their alignment with human annotators in evaluating label flipping remains considerably weak (Figure 2, Appendix G). This contrast highlights that strong downstream task performance does not lead to reliable label flipping evaluation, and it underscores the need for careful judge model selection.
@@
LLMs may struggle to capture nuanced changes when determining the label of imperfect counterfactuals, as the context may not have fully shifted to support a different label, resulting in ambiguity (Figure 1, Appendix F). Notably, even the optimal judge model fails to fully capture label flipping, exhibiting an average discrepancy of 22.78% relative to the user study results (Table 4 in Appendix G), which implies that a fully automated CDA pipeline is insufficient and necessitates human oversight.
@@
We investigate how the choice of relationship affects CDA outcomes (Appendix H). We notice that when the LFR of $LLM_{\mathcal{J}}$, given its relationship to $LLM_{\mathcal{G}}$, aligns more closely with user study outcomes, the labels identified by $LLM_{\mathcal{J}}$ are associated with improved performance and greater robustness on both unseen and out-of-distribution data Kaushik et al. (2020); Gardner et al. (2020) (Appendix H). This association is particularly evident when these identified labels are treated as the ground-truth labels for counterfactuals, which subsequently serve as data points for CDA. This can be attributed to the fact that noisy and incorrect labels provided by judge models with suboptimal relationships contribute to performance deterioration Zhu et al. (2022); Song et al. (2023).
@@ 6 Conclusion
In our experiments, we exclusively use models from Qwen and Llama families to generate counterfactuals, as from DeepSeek-R1 DeepSeek-AI (2025) distilled Qwen2.5 and Llama3 models are officially provided [^7] https://huggingface.co/collections/deepseek-ai/deepseek-r1-678e1e131c0169c0bc89728d and can be used out-of-the-box. Additional work is required when employing models from a different model family as $LLM_{\mathcal{G}}$, including using DeepSeek-R1 to generate synthetic data and fine-tuning $LLM_{\mathcal{G}}$ to derive $LLM_{\mathcal{J}}$ with distillation relationships. Between the model families (Qwen, Llama, Mistral, DeepSeek), there are lots of architectural equivalences and similarities, e.g., the same attention (grouped-query attention), position embeddings (RoPE), normalization (RMSNorm) or FFN activation (SwiGLU). We argue that, based on our comprehensive experiments and large-scale user study ($n=90$), our results are considerably robust and generalizable given the similar architectures compared to other model families.
For label flipping evaluation, we select BERT, RoBERTa, Phi4-14B, Mistral-Large-Instruct-2411 and Gemini-1.5-pro as representatives of independent ($\mathcal{R}_{imw}$, §3.2) LLMs with different parameter sizes, without comprehensively assessing models from all other widely known model families. In particular, we evaluate only open-source models, rather than closed-source, proprietary models.
Beyond LFR, counterfactuals can be evaluated subjectively – via human judgment or LLM-as-a-Judge – along dimensions such as coherence, understandability, feasibility, fairness, and completeness Nguyen et al. (2024); Domnich et al. (2025); Wang et al. (2025b). While automated metrics exist for other aspects (e.g., similarity, diversity), in this paper we focus on identifying which generator–judge relationship is preferable for verifying label flipping, as informed by our user-study results.
@@
Acknowledgment
We thank Leonhard Hennig for his review of an earlier paper draft of our paper. We are indebted to the anonymous reviewers of INLG 2025 for their helpful and rigorous feedback. This work has been supported by the Federal Ministry of Research, Technology and Space (BMFTR) as part of the projects newspolygraph (03RU2U151C), BIFOLD 24B and VERANDA (16KIS2047).
Appendix A Counterfactual Data Augmentation Pipeline
Figure 4 illustrates the CDA pipeline. LLMs first generate counterfactuals using counterfactual generation approaches such as FIZLE and FLARE, both of which are used in our work (§4.1). The generated counterfactuals must then be validated—either by human annotators or judge models—as invalid counterfactuals with incorrect labels may degrade model performance Zhu et al. (2022); Song et al. (2023). Finally, valid counterfactuals are utilized as additional training data to enhance model performance and robustness.
Figure 4: Counterfactual Data Augmentation pipeline.
Appendix B Datasets
B.1 Label Distribution
@@ Appendix C Models
Gemini-1.5-pro Gemini (2024) n.a. https://gemini.google.com/
@@
Appendix D Downstream Task Performance
@@
Appendix E Human Annotation
@@
Appendix F Challenges in Label Flipping Identification
@@
Appendix G Automatic Evaluation
G.1 Average Ranking
Model Relation. AG News SST2 SNLI Avg.
Llama3-8B
$\mathcal{R}_{sm}$ 1 7 4 4
$\mathcal{R}_{dm}$ 3 9 3 5
$\mathcal{R}_{sf}$ 4 8 1 4.33
$\mathcal{R}_{imw}$ 8 3 8 6.33
$\mathcal{R}_{imwo}$ 4.33 3.67 4.33 4.11
Ensemble 8 4 8 6.67
Qwen2.5-14B
$\mathcal{R}_{sm}$ 4 8 2 4.67
$\mathcal{R}_{dm}$ 3 9 3 5
$\mathcal{R}_{sf}$ 2 6 4 4
$\mathcal{R}_{imw}$ 8 2 8 6
$\mathcal{R}_{imwo}$ 4 5.33 4 4.44
Ensemble 8 2 8 6
Qwen2.5-32B
$\mathcal{R}_{sm}$ 6 7 6 6.33
$\mathcal{R}_{dm}$ 1 8 5 4.67
$\mathcal{R}_{sf}$ 7 6 4 5.67
$\mathcal{R}_{imw}$ 6.5 2.5 8.5 5.83
$\mathcal{R}_{imwo}$ 3.33 4.67 4 4
Ensemble 8 5 1 4.67
Llama3-70B
$\mathcal{R}_{sm}$ 3 9 2 4.67
$\mathcal{R}_{dm}$ 4 7 3 4.67
$\mathcal{R}_{sf}$ 2 6 5 4.33
$\mathcal{R}_{imw}$ 8 4.5 8 6.83
$\mathcal{R}_{imwo}$ 4 3.67 6 4.56
Ensemble 8 3 1 4
(a) Counterfactual examples generated using FIZLE.
Model Relation. AG News SST2 SNLI Avg.
Llama3-8B
$\mathcal{R}_{sm}$ 2 9 7 6
$\mathcal{R}_{dm}$ 3 1 8 4
$\mathcal{R}_{sf}$ 6 3 3 4
$\mathcal{R}_{imw}$ 8 5.5 5 6.17
$\mathcal{R}_{imwo}$ 3.33 5.67 5.33 4.78
Ensemble 8 4 1 4.33
Qwen2.5-14B
$\mathcal{R}_{sm}$ 6 8 4 6
$\mathcal{R}_{dm}$ 5 9 8 7.33
$\mathcal{R}_{sf}$ 4 7 7 6
$\mathcal{R}_{imw}$ 8 3.5 3.5 5
$\mathcal{R}_{imwo}$ 2 3.33 5.33 3.55
Ensemble 8 4 3 5
Qwen2.5-32B
$\mathcal{R}_{sm}$ 5 6 4 5
$\mathcal{R}_{dm}$ 4 9 8 7
$\mathcal{R}_{sf}$ 2 8 5 5
$\mathcal{R}_{imw}$ 8.5 2.5 4.5 5.17
$\mathcal{R}_{imwo}$ 3.33 4.33 4 3.89
Ensemble 7 4 7 6
Llama3-70B
$\mathcal{R}_{sm}$ 6 8 4 6
$\mathcal{R}_{dm}$ 4 9 8 7
$\mathcal{R}_{sf}$ 2 2 3 2.33
$\mathcal{R}_{imw}$ 8 6 3 5.67
$\mathcal{R}_{imwo}$ 3 2.67 6 3.89
Ensemble 8 6 6 6.67
(b) Counterfactual examples generated using FLARE.
Table 3: The average ranking of judge-generator model relationship based on $\Delta$ in Table 4 (lower rankings indicate better alignment). Counterfactual examples are generated by Qwen2.5-{14B,32B} and Llama3-{8B,70B}, evaluated across judge models exhibiting same, distilled, same family, and independent w/ and w/o fine-tuning relationships on AG News, SST2 and SNLI. Green-highlighted values indicate that the judge model with the given relationship aligns closely with human annotators, while red-highlighted values indicate the opposite.
Table 3 shows the average ranking of each judge-generator model relationship based on $\Delta$ in Table 4. Figure 3 shows the average ranking and $r_{m/\ell}$ of the relationships $\mathcal{R}_{\mathcal{G},\mathcal{J}}$. We observe that judge models with $\mathcal{R}_{imwo}$ achieve the lowest average ranking, while those with $\mathcal{R}_{imw}$ achieve the highest. Here, a lower ranking indicates that the corresponding label-flipping results are more closely aligned with human evaluation outcomes.
G.2 LFR Differences in Discrete Values
Model Relation. Judge Model ($LLM_{\mathcal{J}}$) AG News SST2 SNLI
Llama3-8B
$\mathcal{R}_{sm}$ Llama3-8B +15.36+15.36 23.53-23.53 +27.00+27.00
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Llama-8B +17.36+17.36 27.13-27.13 +26.80+26.80
$\mathcal{R}_{sf}$ Llama3-70B +22.56+22.56 25.53-25.53 +20.80+20.80
$\mathcal{R}_{imw}$ BERT +36.56+36.56 15.33-15.33 +32.80+32.80
$\mathcal{R}_{imw}$ RoBERTa +37.56+37.56 23.13-23.13 +35.80+35.80
$\mathcal{R}_{imwo}$ Phi4-14B +15.56+15.56 23.53-23.53 +26.80+26.80
$\mathcal{R}_{imwo}$ Mistral-Large +23.56+23.56 21.13-21.13 +32.40+32.40
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +28.32+28.32 17.65-17.65 +30.60+30.60
- Ensemble +37.56+37.56 23.13-23.13 +35.80+35.80
- User Study $75.56$ $46.67$ $100.00$
Qwen2.5-14B
$\mathcal{R}_{sm}$ Qwen2.5-14B +21.24+21.24 28.62-28.62 +17.80+17.80
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Qwen-14B +20.64+20.64 29.02-29.02 +18.40+18.40
$\mathcal{R}_{sf}$ Qwen2.5-72B +17.04+17.04 28.02-28.02 +21.20+21.20
$\mathcal{R}_{imw}$ BERT +46.04+46.04 23.02-23.02 +31.80+31.80
$\mathcal{R}_{imw}$ RoBERTa +43.84+43.84 24.02-24.02 +38.60+38.60
$\mathcal{R}_{imwo}$ Phi4-14B +14.84+14.84 27.62-27.62 +13.80+13.80
$\mathcal{R}_{imwo}$ Mistral-Large +27.04+27.04 28.42-28.42 +22.00+22.00
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +22.13+22.13 26.62-26.62 +27.20+27.20
- Ensemble +43.84+43.84 24.02-24.02 +32.02+32.02
- User Study $84.44$ $57.78$ $100.00$
Qwen2.5-32B
$\mathcal{R}_{sm}$ Qwen2.5-32B 10.33-10.33 22.13-22.13 +30.00+30.00
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Qwen-32B 0.53-0.53 22.33-22.33 +25.60+25.60
$\mathcal{R}_{sf}$ Qwen2.5-72B 16.21-16.21 21.73-21.73 +23.80+23.80
$\mathcal{R}_{imw}$ BERT +9.62+9.62 19.53-19.53 +33.00+33.00
$\mathcal{R}_{imw}$ RoBERTa +27.62+27.62 15.13-15.13 +38.20+38.20
$\mathcal{R}_{imwo}$ Phi4-14B 3.98-3.98 22.93-22.93 +15.80+15.80
$\mathcal{R}_{imwo}$ Mistral-Large +10.22+10.22 18.13-18.13 +22.60+22.60
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +3.40+3.40 18.76-18.76 +30.60+30.60
- Ensemble +27.62+27.62 20.93-20.93 +12.80+12.80
- User Study $62.22$ $66.67$ $100.00$
Llama3-70B
$\mathcal{R}_{sm}$ Llama3-70B +1.84+1.84 43.38-43.38 +16.40+16.40
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Llama-70B +3.84+3.84 41.70-41.70 +18.40+18.40
$\mathcal{R}_{sf}$ Llama3-8B 1.76-1.76 40.58-40.58 +26.20+26.20
$\mathcal{R}_{imw}$ BERT +18.24+18.24 36.98-36.98 +31.40+31.40
$\mathcal{R}_{imw}$ RoBERTa +21.24+21.24 36.78-36.78 +37.40+37.40
$\mathcal{R}_{imwo}$ Phi4-14B +0.84+0.84 42.38-42.38 +22.80+22.80
$\mathcal{R}_{imwo}$ Mistral-Large +11.64+11.64 31.38-31.38 +29.80+29.80
$\mathcal{R}_{imwo}$ Gemini-1.5-pro 10.67-10.67 35.17-35.17 +31.40+31.40
- Ensemble +21.24+21.24 36.78-36.78 +9.80+9.80
- User Study $64.44$ $42.22$ $100.00$
(a) Counterfactual examples generated using FIZLE.
Model Relation. Judge Model ($LLM_{\mathcal{J}}$) AG News SST2 SNLI
Llama3-8B
$\mathcal{R}_{sm}$ Llama3-8B +50.27+50.27 $11.29$ +34.86+34.86
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Llama-8B +54.87+54.87 1.10-1.10 +40.46+40.46
$\mathcal{R}_{sf}$ Llama3-70B +64.47+64.47 +4.29+4.29 +29.26+29.26
$\mathcal{R}_{imw}$ BERT +68.27+68.27 +8.77+8.77 +31.04+31.04
$\mathcal{R}_{imw}$ RoBERTa +68.07+68.07 +6.24+6.24 +29.77+29.77
$\mathcal{R}_{imwo}$ Phi4-14B +56.07+56.07 +3.61+3.61 +28.50+28.50
$\mathcal{R}_{imwo}$ Mistral-Large +58.47+58.47 +9.34+9.34 +62.34+62.34
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +39.43+39.43 +9.01+9.01 +30.60+30.60
- Ensemble +68.07+68.07 +6.24+6.24 +14.25+14.25
- User Study $86.67$ $73.33$ $100.00$
Qwen2.5-14B
$\mathcal{R}_{sm}$ Qwen2.5-14B $46.93$ 22.44-22.44 +32.29+32.29
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Qwen-14B +42.13+42.13 24.16-24.16 +46.71+46.71
$\mathcal{R}_{sf}$ Qwen2.5-72B +41.93+41.93 22.44-22.44 +36.05+36.05
$\mathcal{R}_{imw}$ BERT +53.53+53.53 16.93-16.93 +30.09+30.09
$\mathcal{R}_{imw}$ RoBERTa +53.13+53.13 19.00-19.00 +33.23+33.23
$\mathcal{R}_{imwo}$ Phi4-14B +41.13+41.13 22.44-22.44 +33.23+33.23
$\mathcal{R}_{imwo}$ Mistral-Large +40.73+40.73 12.46-12.46 +54.55+54.55
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +11.02+11.02 17.75-17.75 +27.20+27.20
- Ensemble +53.13+53.13 19.00-19.00 +33.23+33.23
- User Study $73.33$ $66.67$ $100.00$
Qwen2.5-32B
$\mathcal{R}_{sm}$ Qwen2.5-32B +38.31+38.31 35.36-35.36 +32.54+32.54
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Qwen-32B +37.71+37.71 39.49-39.49 +50.36+50.36
$\mathcal{R}_{sf}$ Qwen2.5-72B +35.91+35.91 37.54-37.54 +34.68+34.68
$\mathcal{R}_{imw}$ BERT +47.91+47.91 31.92-31.92 +30.64+30.64
$\mathcal{R}_{imw}$ RoBERTa +47.41+47.41 32.26-32.26 +34.92+34.92
$\mathcal{R}_{imwo}$ Phi4-14B +36.71+36.71 36.39-36.39 +30.64+30.64
$\mathcal{R}_{imwo}$ Mistral-Large +38.91+38.91 24.92-24.92 +53.21+53.21
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +12.29+12.29 34.32-34.32 +30.60+30.60
- Ensemble +47.31+47.31 32.26-32.26 34.92-34.92
- User Study $71.11$ $51.11$ $100.00$
Llama3-70B
$\mathcal{R}_{sm}$ Llama3-70B +39.13+39.13 25.05-25.05 +36.52+36.52
$\mathcal{R}_{dm}$ DeepSeek-R1-Distill-Llama-70B +34.73+34.73 27.11-27.11 +44.58+44.58
$\mathcal{R}_{sf}$ Llama3-8B +26.53+26.53 12.09-12.09 +33.25+33.25
$\mathcal{R}_{imw}$ BERT +42.13+42.13 22.53-22.53 +30.48+30.48
$\mathcal{R}_{imw}$ RoBERTa +39.73+39.73 23.67-23.67 +37.53+37.53
$\mathcal{R}_{imwo}$ Phi4-14B +33.53+33.53 20.69-20.69 +37.78+37.78
$\mathcal{R}_{imwo}$ Mistral-Large +37.53+37.53 11.63-11.63 +47.60+47.60
$\mathcal{R}_{imwo}$ Gemini-1.5-pro +19.56+19.56 15.17-15.17 +31.40+31.40
- Ensemble +39.73+39.73 23.67-23.67 +37.53+37.53
- User Study $73.33$ $62.22$ $100.00$
(b) Counterfactual examples generated using FLARE.
Table 4: The LFR difference ($\Delta\%$) between the user study and the judge–generated model relationships (with values closer to 0 indicating better alignment). Counterfactual examples are generated by Qwen2.5-{14B,32B} and Llama3-{8B,70B}, evaluated across judge models exhibiting same, distilled, same family, and independent w/ and w/o fine-tuning relationships on AG News, SST2 and SNLI. The user study to assess the LFR is conducted on 45 selected counterfactuals (§4.4). Red-highlighted values indicate that the judge model with the given relationship aligns closely with human annotators, while green-highlighted values indicate the opposite.
Table 4 shows the differences calculated by subtracting the user study results from the results of the judge models $LLM_{\mathcal{J}}$ results.
G.3 A Sanity Check